OpenStudy (anonymous):

I cannot line up a fraction equation so it can stay lined up when someone views it. Is there a way to set it so it can be viewed properly.

OpenStudy (amistre64):

\[\frac{top}{bottom}\]

OpenStudy (amistre64):

right click that and select "show source"

OpenStudy (amistre64):

\[\text{\[\frac{top}{bottom}\ ]}\]

OpenStudy (anonymous):

If you don't want to bother with all that just use the divide symbol / 1/2 + 1/2 = 1 or y = 1x/2 + b

OpenStudy (amistre64):

1 3 4 --- + --- = --- = 2 ... good ole ascii :) 2 2 2

OpenStudy (anonymous):

\[\frac{1+\frac{1}{x}}{1-\frac{1}{x}}=\frac{x+1}{x-1}\]

OpenStudy (anonymous):

ascii and ye shall receive

OpenStudy (anonymous):

1 + 1/x / 1 -1/x = x + 1 / x-1 Although I can read it, it is a mess but most people could read it here lol